Express the quadratic function h(x)=(x-r)(x-s) in standard form, and identify a,b, and c.
Answer #1

Given that h(x) = (x-r)(x-s)

i.e., h(x) = x2-(r+s)x+r*s

This is the standard form.

Comparing with h(x) = ax2+bx+c we get,

a = 1, b = -(r+s), c = r*s.